Shift summary for Eve, 7/18/99

Ed Blucher (blucher@hep.uchicago.edu)
Sun, 18 Jul 1999 23:58:29 -0500

Shift Summary for Evening Shift, 07/18/99

On Shift: EB,MM

Hours of beam available at all intensities: 8.4
Hours of beam available at nominal intensity: 8.4
Hours of E832 data (special runs not included): 8.2
Hours of special runs (laser scans, muon runs...): 0

Accounting of lost time: tape change, KFRENDs

========================================================================
Runs ended this shift:
Run 13944 - BEAM_832 - 27440 KL ->2pi0

Tapes brought to FCC this shift: QKC152-154,QKS152-154,QKK152-154,QKB052

========================================================================
Accesses during shift: none

========================================================================
Overview of shift:
E832 Datataking

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
13944 18:30 New ArEth bottles connected. Two new He tubes (3+4)
connected.

13944 22:45 End of run - tapes full.
Note: Two filter processes on ktev2 took ~10% more
cpu time per event throughout run. L3proc brute force
was set at start of next run to fix problem.

13945 22:56 Start run

========================================================================
Follow up Notes on Previous Shift

No problems with CsI 1769

========================================================================
Items to Follow up on Next Shift